# Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits

### v1

@@ III Numerical results
We next apply the proposed ML model to predict properties of the state evolved by a global Hamiltonian $\mathsf{H}=J_{1}\otimes_{i=1}^{N}Z_{i}+\sum_{i=1}^{N}X_{i}$, where $J_{1}$ is a predefined constant and $N$ is the qubit count with $N=60$. The initial state is fixed to be |$\ket{0}^{\otimes N}$ and the circuit $U(\bm{x})$ is the Trotterized time evolution of $\mathsf{H}$. By properly selecting the evolution time at each Trotter step and $J_{1}$, the Trotterized time-evolution circuit takes the form as $U(\bm{x})=\prod_{j=1}^{d}(e^{-\imath\bm{x}_{j}\otimes_{i=1}^{N}Z_{i}}\otimes_{i=1}^{N}\operatorname{RX}(\pi/3))$. We consider two settings of the total number of Trotter steps, i.e., $d=1$ and $d=30$. For both cases, at the training stage, we constitute the dataset $\mathcal{T}_{\mathsf{s}}$ following the same rule presented in the last simulation. The only difference is the dataset size and the shot number, which is $n=20$ and $T=500$ in the former case, and $n=50000$ and $T=50$ in the latter case.
The first subtask is predicting the magnetization with $\braket{\bar{Z}}=\frac{1}{60}\sum_{i}\braket{Z_{i}}$ when $d=1$. The comparison between the prediction with $\Lambda=1$ (full expansion) and the exact value is shown in Fig. 3(a). Namely, we select $25$ test examples evenly distributed across the interval $[-\pi,\pi]$. By increasing the number of training examples $n$ from $10$ to $20$, the prediction values of the proposed ML model almost match the exact results.
The second subtask is much more difficult, which is predicting the magnetization with $\braket{\bar{Z}}$ when $d=30$. In this scenario, full expansion becomes computationally infeasible, where the cardinality of the frequency set is $3^{30}$. In this regard, we evaluate the performance of the proposed ML models by adopting low truncation values with $\Lambda\in\{1,2,3,4\}$. Fig. 3(b) demonstrates the RMS prediction error of our proposal with the varied number of training examples, i.e., $n\in\{10,10^{2},10^{3},10^{4}\}$. An immediate observation is that with the increased $n$, the averaged prediction error continuously decreases for all settings of $\Lambda$. In addition, the achieved results hint that the proposed model with a larger truncation value $\Lambda$ may require more training examples to surpass the one with a smaller truncation value. For instance, the proposed ML model with $\Lambda=2$ attains a better performance compared to $\Lambda=1$ when $n=10^{4}$. It is expected that the model with $\Lambda=3,4$ can attain better performance by further increasing $n$.
IV Discussion and outlook
@@
Acknowledgement
The authors thank Hsin-Yuan Huang, Beng Yee Gan, and Po-Wei Huang for useful discussions.
Supplementary Material of “Efficient Learning for Linear Properties of Bounded-Gate Quantum Circuits”
@@ A.1 Classical shadow and its application in estimating incoherent dynamics of quantum states
where $\mathbb{P}(\bm{b})=\bra{\bm{b}}U\rho U^{\dagger}\ket{\bm{b}}$. Such a linear map implies that the unknown state $\rho$ can be formulated as
@@
In other words, the state $\rho$ can be estimated by sampling the snapshot with $M$ times following the distribution $\mathbb{P}(\bm{b})$. Define the $t$-th snapshot as $U_{t}^{\dagger}\ket{\bm{b}_{t}}\bra{\bm{b}_{t}}U_{t}$ with $t\in[T]$ and $U_{t}\sim\mathcal{U}$. The estimated state of $\rho$ is
$\hat{\rho}_{T}=\frac{1}{T}\sum_{t=1}^{T}\hat{\rho}_{t},\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ \hat{\rho}_{t}=\mathcal{M}^{-1}(U_{t}^{\dagger}\ket{\bm{b}_{t}}\bra{\bm{b}_{t}}U_{t}).$ (12)
@@
$|\cdot\rangle\rangle=\operatorname{Tr}(\cdot P_{i})\leavevmode\nobreak\ \text{with}\leavevmode\nobreak\ P_{i}\in\{\mathbb{I},X,Y,Z\}^{N}.$ (15)
For example, the state |$\ket{0}^{\otimes N}$ satisfies (|$(\ket{0}\bra{0})^{\otimes N}=((\mathbb{I}+Z)/2)^{\otimes N}$, which indicates that its representation under the Pauli basis yields
|$\ket{0}^{\otimes N}\equiv|\bm{0}\rangle\rangle=((\mathbb{I}+Z)/2)^{\otimes N}=\left([1,0,0,1]^{\top}\right)^{\otimes N}.$ (16)
@@
$[\operatorname{\mathfrak{U}}(\bm{\theta})]_{ij}=\langle\langle P_{i}|\operatorname{\mathfrak{U}}(\bm{\theta})|P_{j}\rangle\rangle=\operatorname{Tr}\left(P_{i}U(\bm{\theta})P_{j}U(\bm{\theta})^{\dagger}\right).$ (18)
@@
$\operatorname{\mathfrak{R}_{Z}}(\bm{x}_{j})=\begin{bmatrix}1&0&0&0\\ 0&\cos(\bm{x}_{j})&\sin(\bm{x}_{j})&0\\ 0&-\sin(\bm{x}_{j})&\cos(\bm{x}_{j})&0\\ 0&0&0&1\end{bmatrix}=D_{0}+\cos(\bm{x}_{j})D_{1}+\sin(\bm{x}_{j})D_{2},$ (19)
where $D_{0}=\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$, $D_{1}=\begin{bmatrix}0&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix}$, and $D_{2}=\begin{bmatrix}0&0&0&0\\ 0&0&-1&0\\ 0&1&0&0\\ 0&0&0&0\end{bmatrix}$.
@@
$\rho(\bm{x})=U(\bm{x})(\ket{0}\bra{0})^{\otimes N}U(\bm{x})^{\dagger}=\sum_{\bm{\omega}}\Phi_{\bm{\omega}}(\bm{x})\langle\langle 0|\operatorname{\mathfrak{U}}_{\bm{\omega}}^{\dagger}\equiv\sum_{\bm{\omega}}\Phi_{\bm{\omega}}(\bm{x})\rho_{\bm{\omega}}.$ (20)
@@
$\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\leavevmode\nobreak\ &\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\leavevmode\nobreak\ \bm{\omega}_{i}=-1\end{cases}.$ (21)
@@ A.3 Generality of exploring the learnability of $\mathcal{F}$
=$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}\left(U(\bm{x})\rho_{0}U(\bm{x})^{\dagger}O\right)\Big{|}\bm{x}\in[-\pi,\pi]^{d},O\sim\mathbb{D}_{O}\right\},$ (23)
[…]
@@
where $\eta$ is the learning rate, $c_{1}\in\mathbb{R}$ is the target label, and $f(\bm{x}^{(t)},O)$ amounts to the output of the quantum circuit defined in Eq. (23). The optimization is terminated when the training loss is lower than a threshold or the total number of iterations $T$ exceeds a predefined value. Two main protocols of VQAs are quantum neural networks (QNNs) and variational quantum Eigen-solver (VQEs). The former is utilized to solve machine learning tasks such as classification and regression, and the latter is exploited to estimate the ground state energy of a given Hamiltonian. Notably, when QNNs are applied, the classical input $\bm{x}$ should be divided into two parts, where the first part is used to encode training examples (without updating) and the second part refers to the trainable parameters. Refer to SM J for details.
@@ A.4 Literature review
Our approach distinguishes itself from previous Fourier-based methods by offering broader applicability beyond the classical simulation of VQAs. A specific illustration is that prior Fourier-based approaches often impose constraints on the initial state, such as requiring it to be a simple product state (|$(\ket{0}\bra{0})^{\otimes N}$ interpreted in SM A.2. In contrast, our proposal eliminates this requirement. This advancement is rooted in the hybrid nature of our approach, which utilizes shadow information from quantum computers to build the training dataset. Technically, our method integrates three distinct techniques—classical shadow, kernel method, and trigonometric expansion—paving the way for the development of novel strategies to enhance variational quantum algorithms with provable guarantees.
@@
Table 1: Summary of the key focuses between different works.
@@
Following notations in Eq. (3), denote a dataset $\{\bm{x}^{(i)},\tilde{f}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples with $\bm{x}^{(i)}\in[\pi,\pi]^{d}$ and $\tilde{f}_{T}(\bm{x}^{(i)})$ being the estimation of $f(\bm{x}^{(i)},O)$ using $T$ incoherent measurements. Then, when $T\leq\frac{(1-\epsilon^{2})}{\epsilon}N\log 2$ and $\epsilon\ll 1$, the training data size
@@
is sufficient and necessary to achieve $\mathbb{E}_{\bm{x}\sim\mathbb{D}_{\mathcal{X}}}\left|h(\bm{x},O)-\operatorname{Tr}(\rho(\bm{x}),O)\right|\leq\epsilon$ with high probability. However, under the assumption that $\mathsf{BQP}\not\subset\mathsf{HeurP}^{\mathbb{D}^{\prime}}/\mathsf{poly}$ for some distribution $\mathbb{D}^{\prime}$, there exists a class of $G$-bounded-gate quantum circuits requiring $\Omega(2^{d})$ runtime cost to achieve $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left|h(\bm{x},O)-\operatorname{Tr}(\rho(\bm{x}),O)\right|\leq\epsilon$.
@@
Consider that a learner collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[\pi,\pi]^{d}$. Suppose that with high probability, the learned model $h(\cdot,\cdot)$ can achieve
@@
Consider a learner quires an $N$-qubit quantum circuits and collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[\pi,\pi]^{d}$. Suppose with high probability, there exists a learned model $h(\cdot,\cdot)$ that can achieve
@@
[…]
@@
unless $\mathsf{BQP}\subset\mathsf{HeurP}/\mathsf{Poly}$.
@@
Theorem (Restatement of Theorem  B.1).
Consider that a learner collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[\pi,\pi]^{d}$. Suppose that with high probability, the learned model $h(\cdot,\cdot)$ can achieve
@@
To reach Theorem B.1, we constitute a class of quantum circuits and analyze the lower bound of the required sample complexity for a learner to achieve the average prediction error $\epsilon$. In particular, we consider a simple family of two-qubit quantum circuits illustrated in Fig. C.4(a), which is composed of $d$ $\operatorname{RZ}$ gates and at most $6d$ $\operatorname{CNOT}$ gates, as each SWAP gate can be decomposed into three $\operatorname{CNOT}$ gates. The explicit expression of this class of quantum circuits is
=$\mathcal{F}=\left\{f_{\bm{a}}(\bm{x})=\operatorname{Tr}\left(W(\bm{x},\bm{a})(\rho_{0}\otimes\ket{0}\bra{0})W(\bm{x},\bm{a})^{\dagger}O\right)\big{|}\bm{a}\in\{0,1\}^{d},\bm{x}\sim[-\pi,\pi]^{d}\right\},$ (33)
where $\rho_{0}\otimes\ket{0}\bra{0}$ | is the initial two-qubit state with
@@
Figure C.4: (a) Illustration of the constructed family of quantum circuits. The dashed box refers that $\forall i\in[d]$, the SWAP gate $G_{i}$ is optionally applied to the circuit if the bit string satisfies $\bm{a}_{i}=1$. (b) The visualization of the simplified circuit when $\bm{a}=00...1...0$ (upper) and $\bm{a}=00...0...1$ (lower). The dashed green box stands for replacing the $\operatorname{RZ}$ gate with the identity gate.
Note that any candidate in $\mathcal{F}$ can be simplified to a single-qubit circuit, as demonstrated in Fig. C.4(b). That is, the equivalent single-qubit circuit only consists of a sequence of $\operatorname{RZ}(\bm{x}_{j})$ gates whose index $\bm{a}_{j}$ is zero applied to the initial state $\rho_{0}$, followed by the measurement operator $BX$. Thus, the simplified expression of this family of circuits is
=$\mathcal{F}=\left\{f_{\bm{a}}(\bm{x})=B\operatorname{Tr}\left(V^{\bm{a}_{d}}(\bm{x}_{d})\cdots V^{\bm{a}_{1}}(\bm{x}_{1})\rho_{0}V^{\bm{a}_{1}}(\bm{x}_{1})^{\dagger}\cdots V^{\bm{a}_{d}}(\bm{x}_{d})^{\dagger}X\right)\big{|}\bm{a}\in\{0,1\}^{d},\bm{x}\sim[-\pi,\pi]^{d}\right\},$ (36)

### v2

@@ RESULTS
[…]
The comparison between the exact value and the prediction with $\Lambda=1$ (full expansion) is shown in Fig. 3(a). We select $25$ test examples evenly distributed across the interval $\pi,\pi]$. By increasing the number of training examples $n$ from $10$ to $20$, the prediction values of the proposed ML model almost match the exact results.
[…]
The pretraining process involves three steps. In the first two steps, we construct the training dataset $\mathcal{T}_{\mathsf{s}}$ by randomly sampling $\bm{x}\in[-\pi,\pi]^{99}$ from a uniform distribution and inputting these values into the ansatz $U(\bm{x})$ to generate classical shadows. The number of trainable examples is set as $n=1500$ and the snapshot for each quantum state $\rho(\bm{x})$ has two settings, which are $T=300$ (shadow case) and $T\rightarrow\infty$ (ideal case). Once $\mathcal{T}_{\mathsf{s}}$ is collected, we move to the third step, where the proposed ML model $h_{\mathsf{s}}(\bm{x},\mathsf{H})$ built on $\mathcal{T}_{\mathsf{s}}$ to pretrain VQEs. The pretraining amounts to an optimization task with $\min_{\bm{x}}h_{\mathsf{s}}(\bm{x},\mathsf{H})$. The Adam optimizer is used to minimize this surrogate loss, with a learning rate of $0.01$ and a total of $500$ iterations. Refer to SM K for more details.
The achieved results are illustrated in Fig. 3(b). The inset illustrates the prediction error between the exact result and the optimized $h_{\mathsf{s}}(\bm{x},\mathsf{H})$, evaluated using $500$ test examples sampled from the same data distribution as the training examples. The prediction error for both the shadow and the ideal datasets remains remarkably low, demonstrating the success of the learning process. The main plot shows the dynamics of the proposed ML model during the minimization of the surrogate loss. For both the noisy and ideal cases, the estimated energies after pretraining are close to the exact value.
DISCUSSION
@@
METHODS
Computational efficiency. The proposed ML model, as mentioned in the main text, is not only sample efficient but also computationally efficient when the quantity $C$ is well bounded. Here we briefly discuss why the proposed ML model is computationally efficient and defer the detailed analysis in SM G.
Recall that our proposal consists of two stages: training dataset collection and model inference. As a result, to demonstrate the efficiency of our proposal, it is equivalent to exhibiting the computational efficiency at each stage.
Training dataset collection. This stage amounts to loading the collected training dataset $\mathcal{T}_{\mathsf{s}}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ to the classical memory. For the classical input controls $\{\bm{x}^{(i)}\}_{i=1}^{n}$, $\mathcal{O}(dn)$ computation cost is sufficient. For the classical shadow $\{\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$, the required computational cost is $\mathcal{O}(nNT)$, where $T$ refers to the number of snapshots. According to Theorem 2, the number of training examples $n$ polynomially scales with $N$ and $d$ for an appropriate $C$. Taken together, this stage is computationally efficient when $C$ is well-bounded.
Inference stage. Following the definition of our model in Eq. (157), its inference involves summing over the assessment of each training example $(\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)}))$, and the evaluation of each training example is composed of two components, i.e., the shadow estimation $\operatorname{Tr}(O\tilde{\rho}_{T}(\bm{x}^{(i)}))$ and the kernel function calculation $\kappa_{\Lambda}(\bm{x},\bm{x}^{(i)})$.
As for the Pauli-based shadow estimation, the computation of each $\operatorname{Tr}(O_{i}\tilde{\rho}_{T}(\bm{x}^{(i)}))$ can be completed in $\mathcal{O}(T)$ time after storing the classical shadow in the classical memory. In other words, the computation can be completed in $\mathcal{O}(Tq)$ runtime with $O=\sum_{i=1}^{q}O_{i}$. As for the kernel function calculation, it can be completed in $\mathcal{O}(|\mathfrak{C}(4C/\epsilon)|)$ runtime.
Taken together, the inference stage is also computationally efficient when $C$ is well bounded, where the required computational cost is $\mathcal{O}(n(Tq+|\mathfrak{C}(4C/\epsilon)|))$.
Proof Sketch of Theorem 1. The proof of Theorem 1 is reached by separately analyzing the lower and upper bounds of the sample complexity, and the lower bound of the runtime complexity.
[…]
As for the upper bound, we leverage hypothesis testing to reformulate its derivation as a task of quantifying the cardinality of the packing net for the concept class $\mathcal{F}$. Particularly, the cardinality is determined by all possible arrangements of $d$ $\operatorname{RZ}$ gates and GdG-d Clifford gates across $N$-qubit wires, which is inherently a combinatorial problem. Through analysis, we obtain that the upper bound of the packing number of $\mathcal{F}$ is $\binom{G}{d}\cdot N^{d}\cdot 3^{G-d}\cdot\binom{N}{2}^{G-d}$. Then, according to the result of statistical learning theory, the required sample complexity logarithmically depends on this cardinality, leading to the linear scaling with $d$ and $G$.
Runtime complexity. To demonstrate the computational hardness, we leverage results from the $\mathsf{BQP}$ complexity class. Specifically, we construct a tailored concept class $\mathcal{F}$ that incorporates a concept $f^{*}$ realized by the $\mathsf{BQP}$ circuit. Consequently, if an efficient classical learner were to exist for this task, it would imply that any $\mathsf{BQP}$ language could be decided in $\mathsf{P/poly}$—a scenario that is widely considered implausible.
[…]
The remaining step to complete the proof involves demonstrating that the quantum state $\rho^{*}(\bm{x})$ can be efficiently prepared using RZ and Clifford gates. Taken together, this completes the proof.
Proof Sketch of Theorem 2. We first exhibit that the expected risk is upper bounded by the truncation error $\text{Err}_{1}$ and the estimation error $\text{Err}_{2}$ with $\mathbb{E}_{\bm{x}}|h_{\mathsf{s}}(\bm{x},O)-f(\rho(\bm{x}),O)|^{2}\leq(\sqrt{\text{Err}_{1}}+\sqrt{\text{Err}_{2}})$, where $\text{Err}_{1}=\mathbb{E}_{\bm{x}}[|\operatorname{Tr}\left(O\rho_{\Lambda}(\bm{x})\right)-\operatorname{Tr}(O\rho(\bm{x}))|^{2}]$ $\text{Err}_{2}=\mathbb{E}_{\bm{x}}[|\operatorname{Tr}\left(O\rho_{\Lambda}(\bm{x})\right)-h_{\mathsf{s}}(\bm{x},O)|^{2}]$. Here we expand the explored state under the trigonometric monomial basis in Eq. (8) with $\rho_{\Lambda}=\sum_{\bm{\omega}\in\mathfrak{C}(\Lambda)}\Phi_{\bm{\omega}}(\bm{x})$ and $\rho=\rho_{d}$. Intuitively, $\text{Err}_{1}$ measures the discrepancy between the target state and the truncated state, and $\text{Err}_{2}$ measures the difference between the prediction and the estimation of the truncated state.
The upper bound of $\text{Err}_{1}$ is established by connecting with the gradient norm of the expectation value $\operatorname{Tr}(\rho(\bm{x})O)$. More concisely, supported by the results of Fourier analysis, we prove that $\text{Err}_{1}$ is upper bounded by $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}/{\Lambda}$. Then, under the assumption $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\|\nabla_{\bm{x}}\operatorname{Tr}(\rho(\bm{x})O)\|_{2}^{2}\leq C$, we obtain $\text{Err}_{1}\leq C/\Lambda$. The upper bound of $\text{Err}_{2}$ is obtained by exploiting the results of Pauli-based shadow estimation. Particularly, we first prove that $h_{\mathsf{s}}(\bm{x},O)$ is an unbiased estimator of $\rho_{\Lambda}(\bm{x})$. Then we make use of the results of Pauli-based shadow to quantify the estimation error, i.e., $\text{Err}_{2}\leq\tilde{\mathcal{O}}(|\mathfrak{C}(\Lambda)|\frac{1}{2n}B^{2}9^{K})$. Taken together, we reach the complexity bound in Theorem 2.
Code implementation and simulation details. All numerical simulations in this work are conducted using Python, primarily leveraging the Pennylane and PastaQ libraries. For small qubit sizes with $N\leq 20$, we use Pennylane to generate training examples and make predictions for new inputs. For larger qubit sizes with $N>20$, we utilize PastaQ to generate training data and Numpy to handle post-processing and new input predictions.
In the task of predicting the two-point correlation of rotational GHZ states (i.e., Fig. 2(a)), the test value used in the prediction stage is $\bm{x}=[1.234,-1.344,-1.716]$. All diagonal elements in the correlation matrix are set as $1$.
DATA AVAILABILITY
The data generated and analyzed during the current study are publicly available at the following repositories:
Github: https://github.com/yuxuan-du/Efficient_Predicting_Bounded_Gate_QC;
Google Drive: https://drive.google.com/drive/folders/1lzDyZ7d3Wsm2ncFMGhkUjJg7HrI3ouXy.
These datasets support the findings of the study and include all data necessary to reproduce the results presented in the main text and supplementary materials.
CODE AVAILABILITY
The code used in this study is available at the Github repository https://github.com/yuxuan-du/Efficient_Predicting_Bounded_Gate_QC.
ACKNOWLEDGEMENTS
D.T. is supported by the National Research Foundation, Singapore, under its NRF Professorship Award No. NRF-P2024-001.
AUTHOR CONTRIBUTIONS
The project was conceived by Y.D., and D.T. Theoretical results were proved by Y.D., and M.H. Numerical simulations and analysis were performed by Y.D. All authors contributed to the write-up.
COMPETING INTERESTING
The authors declare no competing interests.
@@ A.1 Classical shadow and its application in estimating incoherent dynamics of quantum states
[…]
@@
$|\cdot\rangle\rangle=\operatorname{Tr}(\cdot P_{i})\penalty 10000\ \text{with}\penalty 10000\ P_{i}\in\{\mathbb{I},X,Y,Z\}^{N}.$ (15)
[…]
@@
$\operatorname{\mathfrak{U}}(\bm{\theta})]_{ij}=\langle\langle P_{i}|\operatorname{\mathfrak{U}}(\bm{\theta})|P_{j}\rangle\rangle=\operatorname{Tr}\left(P_{i}U(\bm{\theta})P_{j}U(\bm{\theta})^{\dagger}\right).$ (18)
@@
$\operatorname{\mathfrak{R}_{Z}}(\bm{x}_{j})=\begin{bmatrix}1&0&0&0\\ 0&\cos(\bm{x}_{j})&\sin(\bm{x}_{j})&0\\ 0&-\sin(\bm{x}_{j})&\cos(\bm{x}_{j})&0\\ 0&0&0&1\end{bmatrix}=D_{0}+\cos(\bm{x}_{j})D_{1}+\sin(\bm{x}_{j})D_{-1},$ (19)
where $D_{0}=\begin{bmatrix}1&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&1\end{bmatrix}$, $D_{1}=\begin{bmatrix}0&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&0\end{bmatrix}$, and $D_{-1}=\begin{bmatrix}0&0&0&0\\ 0&0&-1&0\\ 0&1&0&0\\ 0&0&0&0\end{bmatrix}$.
@@
We now recap the mechanism of LOWESA. When an $N$-qubit quantum circuit $U(\bm{x})$ is composed of $d$ $\operatorname{RZ}$ gates and GdG-d $\operatorname{\text{CI}}$ gates, the state representation under Pauli-basis expansion yields
@@
$\Phi_{\bm{\omega}}(\bm{x})=\prod_{i=1}^{d}\begin{cases}1\penalty 10000\ &\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=0\\ \cos(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=1\\ \sin(\bm{x}_{i})&\textnormal{if}\penalty 10000\ \bm{\omega}_{i}=-1\end{cases}.$ (21)
@@ A.3 Generality of exploring the learnability of $\mathcal{F}$
$\mathcal{F}=\left\{f(\bm{x},O)=\operatorname{Tr}\left(U(\bm{x})\rho_{0}U(\bm{x})^{\dagger}O\right)\Big|\bm{x}\in[-\pi,\pi]^{d}\right\},$ (23)
where $U(\bm{x})=\prod_{l=1}^{d}(\operatorname{RZ}(\bm{x}_{l})u_{e})$ is composed of $d$ $\operatorname{RZ}(\bm{x}_{l})$ gates, GdG-d $\operatorname{\text{CI}}$ gates denoted by $U_{e}$ with $\operatorname{\text{CI}}=\{H,S,\operatorname{CNOT}\}$, $O$ is constituted by multiple local observables with a bounded norm, i.e., $O=\sum_{i=1}^{q}O_{i}$ and $\sum_{l}\|O_{i}\|_{\infty}\leq B$, and the maximum locality of $\{O_{i}\}$ is $K$. This formalism encompasses many applications for quantum computing, including variational quantum algorithms, classical-shadow-based algorithms, and quantum system certification. A common feature of these applications is that the exploited quantum circuit can be hardware-efficient. As a result, for a specified quantum device, its executable quantum circuits can be described by $\operatorname{RZ}+\operatorname{\text{CI}}$ gates with a fixed layout but with different angles, i.e., $U(\bm{x})$ in Eq. (23). Given an unknown state $\rho_{0}$ evolved by $U(\bm{x})$ and incoherently measured by an observable $O$, it forms the mean-value space $\mathcal{F}$ explored in this work. In the following, we detail how $\mathcal{F}$ relates to variational quantum algorithms, classical-shadow-based applications, and quantum system certification, respectively.
@@
where $\eta$ is the learning rate, $c_{1}\in\mathbb{R}$ is the target label, and $f(\bm{x}^{(t)},O)$ amounts to the output of the quantum circuit defined in Eq. (23). The optimization is terminated when the training loss is lower than a threshold or the total number of iterations $T$ exceeds a predefined value [51, 52, 77]. Two main protocols of VQAs are quantum neural networks (QNNs) and variational quantum Eigen-solver (VQEs). The former is utilized to solve machine learning tasks such as classification and regression, and the latter is exploited to estimate the ground state energy of a given Hamiltonian. Notably, when QNNs are applied, the classical input $\bm{x}$ should be divided into two parts, where the first part is used to encode training examples (without updating) and the second part refers to the trainable parameters. Refer to SM K for details.
@@ A.4 Literature review
[…]
@@
Supplementary Table 1: Summary of the key focuses between different works.
@@
Following notations in Eq. (3), denote a dataset $\{\bm{x}^{(i)},\tilde{f}_{T}(\bm{x}^{(i)},O)\}_{i=1}^{n}$ containing $n$ training examples with $\bm{x}^{(i)}\in[-\pi,\pi]^{d}$ and $\tilde{f}_{T}(\bm{x}^{(i)},O)$ being the estimation of $f(\bm{x}^{(i)},O)$ using $T$ incoherent measurements. Then, when $T\leq\frac{(1-\epsilon^{2})}{\epsilon}N\log 2$ and $\epsilon\ll 1$, the training data size
@@
is sufficient and necessary to achieve $\mathbb{E}_{\bm{x}\sim\mathbb{D}_{\mathcal{X}}}\left|h(\bm{x},O)-\operatorname{Tr}(\rho(\bm{x}),O)\right|\leq\epsilon$ with high probability. However, unless $\mathsf{BQP}\subseteq\mathsf{P/poly}$, there exists a class of $G$-bounded-gate quantum circuits that no algorithm can achieve $\mathbb{E}_{\bm{x}\sim[-\pi,\pi]^{d}}\left|h(\bm{x},O)-\operatorname{Tr}(\rho(\bm{x}),O)\right|\leq\epsilon$ in a polynomial time.
@@
Consider that a learner collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[-\pi,\pi]^{d}$. Suppose that with high probability, the learned model $h(\cdot,\cdot)$ can achieve
@@
Consider a learner quires an $N$-qubit quantum circuits and collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[-\pi,\pi]^{d}$. Suppose with high probability, there exists a learned model $h(\cdot,\cdot)$ that can achieve
@@
There exists a class of bounded-gate quantum circuits composed of $d$ $\operatorname{RZ}$ gates and GdG-d $\operatorname{\text{CI}}$ gates such that no polynomial-time algorithm exists to achieve
@@
unless $\mathsf{BQP}\subseteq\mathsf{P/poly}$.
@@
Theorem (Restatement of Theorem B.1).
Consider that a learner collects a dataset $\mathcal{T}=\{\bm{x}^{(i)},\tilde{\rho}_{T}(\bm{x}^{(i)})\}_{i=1}^{n}$ containing $n$ training examples to predict unseen states $\rho(\bm{x})$ with $\bm{x}\in[-\pi,\pi]^{d}$. Suppose that with high probability, the learned model $h(\cdot,\cdot)$ can achieve
@@
To reach Theorem B.1, we constitute a class of quantum circuits and analyze the lower bound of the required sample complexity for a learner to achieve the average prediction error $\epsilon$. In particular, we consider a simple family of two-qubit quantum circuits illustrated in Fig. 1(a), which is composed of $d$ $\operatorname{RZ}$ gates and at most $6d$ $\operatorname{CNOT}$ gates, as each SWAP gate can be decomposed into three $\operatorname{CNOT}$ gates. The explicit expression of this class of quantum circuits is
[…]
where $|\rho_{0}\otimes\ket{0}\bra{0}$ is the initial two-qubit state with
@@
Supplementary Fig. 1: Some intuitions about the proof of Theorem B.1. (a) Illustration of the constructed family of quantum circuits. The dashed box refers that $\forall i\in[d]$, the SWAP gate $G_{i}$ is optionally applied to the circuit if the bit string satisfies $\bm{a}_{i}=1$. (b) The visualization of the simplified circuit when $\bm{a}=00...1...0$ (upper) and $\bm{a}=00...0...1$ (lower). The dashed green box stands for replacing the $\operatorname{RZ}$ gate with the identity gate.
Note that any candidate in $\mathcal{F}$ can be simplified to a single-qubit circuit, as demonstrated in Fig. 1(b). That is, the equivalent single-qubit circuit only consists of a sequence of $\operatorname{RZ}(\bm{x}_{j})$ gates whose index $\bm{a}_{j}$ is zero applied to the initial state $\rho_{0}$, followed by the measurement operator $BX$. Thus, the simplified expression of this family of circuits is
$\mathcal{F}=\left\{f_{\bm{a}}(\bm{x})=B\operatorname{Tr}\left(V^{\bm{a}_{d}}(\bm{x}_{d})\cdots V^{\bm{a}_{1}}(\bm{x}_{1})\rho_{0}V^{\bm{a}_{1}}(\bm{x}_{1})^{\dagger}\cdots V^{\bm{a}_{d}}(\bm{x}_{d})^{\dagger}X\right)\big|\bm{a}\in\{0,1\}^{d},\bm{x}\sim[-\pi,\pi]^{d}\right\},$ (36)